OpenStudy (anonymous):

PLZ HELP WILL GIVE MEDAL AND FAN Make tables to find the solution to 4–x = 2x + 3. Take the integer values of x between −3 and 3.

OpenStudy (thadds2003):

@mathmale

OpenStudy (thadds2003):

@AravindG @mathstudent55 @KendrickLamar2014 @sleepyjess

OpenStudy (anonymous):

this is only part b if you want the whole problem let me know

OpenStudy (thadds2003):

Is this a test?

OpenStudy (anonymous):

no a study guide

OpenStudy (thadds2003):

Ok.

OpenStudy (thadds2003):

What grade are you in? I might be able to solve this myself......

OpenStudy (anonymous):

9th

OpenStudy (thadds2003):

Oh. Let me give it a try...... So this is an inequality, am I correct? If so, x would equal 1.

OpenStudy (anonymous):

yes

OpenStudy (foolofatook):

start by writing the equation down somewhere. then plug in the integers -3 through 3 and eventually u will hac the answer

OpenStudy (thadds2003):

Then you could create the tables........

OpenStudy (anonymous):

how?

OpenStudy (foolofatook):

write down 4–x = 2x + 3. then fill in x for -3 if that dosnt work go to -2 and so on until you get to three

OpenStudy (thadds2003):

Because if I plug it in, 4-1 = 3, and 2*1 -3 =..... -1. Oops, made an error, let me think it through.

OpenStudy (anonymous):

can i just have the graph so i can see how its made and examine it?